OpenStudy (anonymous):

Help me with this radical equation:√(x+10)=-6

OpenStudy (anonymous):

How could you get rid of the square root?

OpenStudy (anonymous):

Square both sides then solve for x. x = 26.

jhonyy9 (jhonyy9):

squared both sides and will get x+10=36 x=36-10 x=26